OpenStudy (anonymous):

I need help solving this problem please explain steps thanks! :) 3x+6/8=7x-1/6

OpenStudy (anonymous):

3x+6/8=7x-1/6 We simplify the equation to the form, which is simple to understand 3x+6/8=7x-1/6 Simplifying: 3x + 0.75=7x-1/6 Simplifying: 3x + 0.75=7x-0.166666666667 We move all terms containing x to the left and all other terms to the right. +3x-7x=-0.166666666667-0.75 We simplify left and right side of the equation. -4x=-0.916666666667 We divide both sides of the equation by -4 to get x. x=0.229166666667

OpenStudy (anonymous):

thank you! :)